A buffer that contains 0.550 M base, B, and 0.229 M of its conjugate acid, BH+, has a pH of 8.90. What is the pH after 0.0020 mol of HCl is added to 0.250 L of this solution?
Answer #1

pott_ 14-pH = 14-8.9= 5.1 [Btt potaplab tog [B] 0-229 0.229 = 5.4p 5.12 pkb tog oss pkb = 5.1 - Log ons [B] = 0.55m pkb = 5.4